OpenStudy (anonymous):

how to prove that a rectangular box with square base and top having a fixed surface area and maximum volume is a cube

OpenStudy (anonymous):

This is an optimization problem

OpenStudy (anonymous):

Do you know how to tackle those?

OpenStudy (anonymous):

if is a cube the V= x^3, A= 2X^2 + 4X^2, derivative A'= 12X so 12X=0, X=0

OpenStudy (anonymous):

Ok, the question says that the S.A. is fixed, correct? So let's give it a S.A. of 500 a "square-based rectangular box" has length the same as width (to keep the base square). This gives you these equations, with w (width and length) and h (height) of the box: surface area (square top/bottom, plus four sides): a = 2w^2 + 4wh volume: v = w^2h Do you follow that?

OpenStudy (anonymous):

yes, i was x's and u r using w,h

OpenStudy (anonymous):



OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Here is the picture to show you visually what I was saying. Now we have our two equations, lets plug in what we know \[\Large 500=2w^2+4wh\] & \[\Large A=w^2h\]

OpenStudy (anonymous):

We are trying to MAXimize Area, so we have to find the max of that function (we'll use derivative in a moment to find it). But Area is in terms of w and h, we need to get it into terms of just w or just h, which one do you want to get it in?

OpenStudy (anonymous):

V= W^2h right?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

We have to get it in terms of one variable, we have the Surface area equation to help us so we have to 1) Isolate either h or w in the surface area equation. 2) Plug it into the area equation So do you want to isolate w or h?

OpenStudy (anonymous):

h= v/w^2

OpenStudy (anonymous):

I advise we go along with h, because I solved it isolating both w and h and h was easier

OpenStudy (anonymous):

We have to isolate w or h in the SURFACE AREA equation and plug it into the **volume (I said area before sorry) formula

OpenStudy (anonymous):

got it, what is the next step?

OpenStudy (anonymous):

We isolate it, can you tell me what you got?

OpenStudy (anonymous):

i isolated h and plugged in V

OpenStudy (anonymous):

Ok, can you tell me what your new Volume formula is?

OpenStudy (anonymous):

V=w^2*(500-2W^2/4w)

OpenStudy (anonymous):

Alright, I'll clean it up for you to \[\Large V= {500w-2w^3 \over 4}\]

OpenStudy (anonymous):

Now we have to take the derivative of that

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Tell me what you get

OpenStudy (anonymous):

v'=(200-24w^2)/16

OpenStudy (anonymous):

2000

OpenStudy (anonymous):

remember that we can factor out the 1/4 when we take the drivative and then we multiply our answer by 1/4 after \[\Large {dy \over dx}[c*f(x)]=c*f \prime x\]

OpenStudy (anonymous):

So it's really the derivative of \[\Large 500w-2w^3\] and then take the answer and put it over 4

OpenStudy (anonymous):

\[\Large {500-6w^2 \over 4}\]

OpenStudy (anonymous):

what about this 250 - 3w^2 / 2

OpenStudy (anonymous):

Tell me how you came up with that result and maybe I can help you with what you're doing wrong

OpenStudy (anonymous):

i just simplified by 2

OpenStudy (anonymous):

Oh, lol I didn't see that, yes those two are the same thing

OpenStudy (anonymous):

\[\Large {250-3w^2 \over2}=0\]

OpenStudy (anonymous):

Solve for x (leave it in radical form)

OpenStudy (anonymous):

w*******

OpenStudy (anonymous):

w=250/3^1/3

OpenStudy (anonymous):

\[\sqrt[3]{250/3}\]

OpenStudy (anonymous):

Ok, it's \[\Large \sqrt{{250 \over 3}}\] sqrt not cuberoot. Now we're going to plug in a point less than that into our derivative equation and then plug in a number greater than that inot our derivative equation to see a max

OpenStudy (anonymous):

point less such as \[\sqrt{250/4}\]